Pappys Potato has come up with a new product, the Potato Pet (they are freeze-dried to last longer). Pappys paid $165,000 f calculate payback period and IRR
Answer #1
Time line 0
Cost of new machine -940000
=Initial Investment outlay -940000
Time line 0 1 2 3 4
Cost of new machine -940000
=Initial Investment outlay -940000
100.00%
Sales 880000 880000 880000 880000
Profits Sales-variable cost 721600 721600 721600 721600
Fixed cost -222000 -222000 -222000 -222000
-Depreciation Cost of equipment/no. of years -235000 -235000 -235000 -235000 0 =Salvage Value
=Pretax cash flows 264600 264600 264600 264600
-taxes =(Pretax cash flows)*(1-tax) 201096 201096 201096 201096
+Depreciation 235000 235000 235000 235000
=after tax operating cash flow 436096.00 436096.00 436096 436096
+Tax shield on salvage book value =Salvage value * tax rate 0
=Terminal year after tax cash flows 0
Total Cash flow for the period -940000 436096.00 436096.00 436096.000 436096
Project
Year Cash flow stream Cumulative cash flow
0 -940000 -940000
1 436096 -503904
2 436096 -67808
3 436096 368288
4 436096 804384
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 2 and 3
therefore by interpolation payback period = 2 + (0-(-67808))/(368288-(-67808))
2.16 Years
Project
IRR is the rate at which NPV =0
IRR 0.302974796
Year 0 1 2 3 4
Cash flow stream -940000 436096 436096 436096 436096
Discounting factor 1 1.302975 1.697743 2.212117 2.8823324
Discounted cash flows project -94000000.00% 334692.6 256868 197139.7 151299.69
NPV = Sum of discounted cash flows
NPV Project = 2.19402E-05
Where
Discounting factor = (1 + IRR)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
IRR= 30.30%